OpenStudy (anonymous):

Find the inverse function for f(x) = ln square root of (x) divided by 5

OpenStudy (loser66):

\[f(x) = \frac{ \ln \sqrt{x} }{ 5 } or f(x)= \ln (\frac{ \sqrt{x} }{ 5 })\]

OpenStudy (loser66):

which one?

OpenStudy (anonymous):

The first one

OpenStudy (loser66):

do you know how to find the inverse function?

OpenStudy (anonymous):

no

OpenStudy (loser66):

let f(x) =y , from now on, I just work with that notation, y = ... first step: switch 2 variables. I mean whenever you see y, replace it by x and vice versa.

OpenStudy (loser66):

you do by yourself, I will correct if you have mistake,

OpenStudy (anonymous):

y=ln square root of (y) divided by 5

OpenStudy (loser66):

that is x =...., not y =...redo, please

OpenStudy (anonymous):

x=ln square root of (y) divided by 5

OpenStudy (loser66):

now solve for that y. i mean you isolate y

OpenStudy (anonymous):

is the answer e^10x^2=y

OpenStudy (loser66):

it's y = e^10x